OpenStudy (anonymous):

\[(\tan ^{2}x-1)\sec x, if cosx= -2/3 and x \in Q2\]

OpenStudy (pratyush5):

Do you have the answer ?

OpenStudy (anonymous):

what is the question??

OpenStudy (anonymous):

what is Q2???

OpenStudy (pratyush5):

Do we have to find value of this = (tan^2x−1)secx ?

OpenStudy (anonymous):

\[(\tan ^{2}x-1)\sec x, if cosx= -2/3 and x \in Q2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont have the answer yet

OpenStudy (pratyush5):